Write an equation of the line that passes through the given point and in perpendicular the the given line. Your answer should be written in slope-intercept form.
P(2,-7),y-2x=-3
Thanks in advance!

Answer by fractalier(6550)

y-2x=-3 means y = 2x - 3.
If the line you want is perpendicular to y-2x=-3 , then the line you want has a slope of (-1/2).
Now plug into y = mx + b to find b...
-7 = (-1/2)(2) + b
-7 = - 1 + b
b = -6
so that
y = (-1/2)x - 6
is the line you want.